OpenStudy (anonymous):

If xy = 8, dx/dt = 2, find dy/dt when x = 3. The more detailed the explanation the better. Thanks in advance!!!

OpenStudy (amistre64):

Dt(xy) = x'y + xy' xy = 8 implies y = 8/x if i see it right hence y' = -8/x^2

OpenStudy (amistre64):

im sure theres an idea brewing in this someplace :)

OpenStudy (amistre64):

hence y' = -8x'/x^2 forgot that x wasnt a fucntion of itself

OpenStudy (anonymous):

I think hamidaarab has come up with something

OpenStudy (amistre64):

then x=3; y = 8/3 if we are to keep the .. i spose he diesnt know that the posts are meant to be responded to inside of them ...

OpenStudy (phi):

xy = 8, dx/dt = 2, find dy/dt when x = 3. picking up amistre's x'y + xy'= 0 sub in x'= 2, x=3 , y= 8/x = 8/3 2*8/3 + 3 y'= 0 solve for y'

OpenStudy (anonymous):

thanks!